# Non-abelian Rees construction and pure motives

### v1

@@ Introduction.
In the most basic case, it relates the tensor category ${\rm{FilVec}}_{K}$ of filtered $K$-vector spaces $(V,F^{\bullet})$ to the tensor category ${\rm{GrVec}}_{K}$ of graded $K$-vector spaces: the map $(V,F^{\bullet})\mapsto\sum x^{-i}K[x]\otimes F^{i}V,$ (a $K[x]$-submodule of $K[x,\frac{1}{x}]\otimes V$) yields a $\mathbb{G}_{m}$-equivariant vector bundle on $\mathbb{A}^{1}={\rm{Spec}}\,K[x]$, and gives rise to a commutative diagram of tensor functors (cf. [8, §3])
@@
The present paper explores the non-abelian generalization of this situation, when the multiplicative group $\mathbb{G}_{m}$ is replaced by an arbitrary reductive $K$-group $G$. This raises immediately two complementary questions:
@@
Question 1 was given a simple and general answer by P. O’Sullivan [36]: roughly speaking, any even $K$-tensor category $\mathcal{C}$} (i.e. such that that any object is killed by some exterior power) is a category of equivariant vector bundles with respect to the tannakian group $G$ of its quotient $\bar{\mathcal{C}}$ by the maximal tensor ideal, on a certain fix-pointed affine $G$-scheme $\rm X$.
This question then becomes: how to describe $\rm X$, and how does $\rm X$ behave with respect to tensor functors ${\mathcal{C}}\to{\mathcal{C}}^{\prime}$ between even categories?
@@
In this paper, we first discuss O’Sullivan’s construction and its functorial properties in full generality (Theorem 1.4.2), and then study in detail the case when $\mathcal{C}$} admits a faithful tensor functor with values in $K$-vector spaces. In this case, under mild extra conditions, it turns out that $\rm X$ is a quasi-homogeneous variety, i.e. a partial compactification of a homogeneous $G$-variety, which is amenable to computation (Theorem 2.3.1).
This gives rise to a Galois-type anti-equivalence between even categories $\mathcal{C}$} and spaces $\rm X$, which one can exploit using the rich toolbox of the theory of affine quasi-homogeneous varieties (Theorem 3.2.1).
@@
The non-abelian Rees construction applies notably to the setting of pure motives, especially motives modulo homological equivalence - modulo some adjustments111to cope with the Künneth projectors, see below.. The standard conjectures predict that such categories $\mathcal{C}$} of homological motives are abelian semisimple (Grothendieck’s standard conjecture D).
@@
In some special instances, the theory of quasi-homogeneous varieties allows to conclude that the deformation space $\rm X$ must be a point, hence that ${\mathcal{C}}=\bar{\mathcal{C}}$ is abelian semisimple. In this way, we get a new proof, and a generalization, of the Clozel-Deligne theorem on numerical equivalence on abelian varieties over finite fields (Theorem 4.5.2).
@@ 1.1. The classical Rees construction and beyond
We are looking for extensions, when $\mathbb{G}_{m}$ is replaced by an arbitrary reductive $K$-group $G$. As mentioned above, this raises two questions Q1, Q2. In order to say a little more about Klyashko-Asok’s partial answer to question Q2, we recall three classical definitions (under the assumption $K=\bar{K}$).
Quasi-homogeneous222a synonym is ‘prehomogeneous’, cf. [8]. $G$-variety.: a $G$-variety $\rm{X}$} which admits a homogeneous space as an open dense subvariety $\rm{X}_{0}$} (informally: a partial compactification of a homogeneous space). We shall be especially concerned with affine embeddings $\rm X$ of a homogeneous space ${\rm X}_{0}=G/H$, cf [7] for a panorama.
Spherical $G$-variety: a quasi-homogeneous $G$-variety which remains quasi-homogeneous if one restricts the action to some Borel subgroup $B$. This condition only depends on the homogeneous space ${\rm X}_{0}$, and can be reformulated as follows: for any irreducible $G$-representation $W,\,dim\,W^{H}\leq 1$ (cf. e.g. [12]). Toric varieties are special cases, cf. e.g. [16].
Fix-pointed $G$-variety: a $G$-variety with a unique $G$-closed orbit, which is a point.
The Brion-Luna-Vust theory provides a combinatorial description of normal spherical varieties (up to codimension $2$ subvarieties) in terms of $1$-parameter subgroups $\mu_{j}:\mathbb{G}_{m}\to G$ and other data. Using this, Asok [8, 5.7] (after Klyachko [26] in the toric case) shows that for any fix-pointed affine normal spherical $G$-variety $\rm{X}$ with open orbit $G/H$, ${\rm{Vec}}_{G}({\rm X})$ is equivalent to a tensor category whose objects are $G$-representations and whose morphisms are $H$-linear maps preserving specific filtrations attached to $\mu_{j}$.
@@ 1.1.3. Question
It is known that every quasi-affine homogeneous variety $\rm X_{0}$ over $K=\bar{K}$ either is affine and does not admit any non-trivial equivariant affine embedding, or admits an equivariant affine embedding $\rm X$ with a fixed point $o$ [7, 3.4]; but it is not clear whether $o$ may be taken to be the unique closed orbit. How restrictive is the existence of a fix-pointed equivariant affine embedding of $\rm X_{0}$?
@@ 1.2. O’Sullivan’s construction
This construction provides a general answer to Q1.
@@
A $K$-tensor category $\mathcal{C}$} is a $K$-linear category333we follow O’Sullivan terminology and do not assume that $\mathcal{C}$} is abelian. equipped with a structure of symmetric monoidal category for which the tensor product is bilinear. The endomorphism ring ${{\mathcal{C}}}({\bf 1},{\bf 1})$ of the unit object is then a commutative $K$-algebra. We say that $\mathcal{C}$} is rigid if every object is dualizable.
@@
This is compatible with the usual terminology for tannakian categories, because if $\mathcal{C}$} is tannakian over $K$, any faithful tensor functor ${\mathcal{C}}\to{\rm{Vec}}_{K^{\prime}}$ is exact [38, 10.7][15, 2.4.1][23]444this result, obtained in different ways, seems to fill the gap in Saavedra’s description of tannakian categories [43] in a more economic way than Deligne’s geometry in tannakian categories [17]..
@@ 1.2.2. Examples and notations
For a quasi-affine $G$-scheme $\rm X$, we denote by ${\rm{Vec}}_{G}({\rm{X}})$ the $K$-tensor category of $G$-equivariant vector bundles on $\rm X$.
@@
A rigid $K$-tensor category $\mathcal{C}$} with ${{\mathcal{C}}}({\bf 1},{\bf 1})=K$ is even if it is pseudo-abelian555i.e. every idempotent endomorphism has a kernel. and every object is annihilated by some exterior power. Then the quotient category
@@
Moreover, $\pi$ admits a (essentially unique) tensor section $\sigma$ [36] (the semisimplicity of $\bar{\mathcal{C}}$ is essential here). By the general criteria of SGA IV, I, 8.11, $\sigma$ then has a (essentially unique) right Ind-adjoint
$\theta:{\rm{Ind}}\,{\mathcal{C}}\to{\rm{Ind}}\,\bar{\mathcal{C}}$
which is a lax symmetric monoidal functor, and $\mathcal{O}({\rm{X}}):=\theta(\bf 1)$ is a commutative algebra in ${\rm{Ind}}\,\bar{\mathcal{C}}$ [36]666cf. also $\rm{nLab}$: monoidal adjunction..
1.2.4.
Let us now assume that $\bar{\mathcal{C}}$ is neutral, and fix a fiber functor
@@
Then $G:=Aut^{\otimes}\bar{\omega}$ is proreductive, $\bar{\mathcal{C}}\stackrel{{\scriptstyle\cong}}{{\to}}{\rm{Rep}}\,G$, and
@@
Here fix-pointed means that there is a $G$-fixed $K$-point $o$, corresponding to a $G$-map $\mathcal{O}({\rm{X}})\to K$ whose restriction to ${\mathcal{O}}({\rm X})^{G}$ is an isomorphism. The $G$-fixed point $o$ is then unique (in ${\rm{REP}}\,G$, ${\mathcal{O}}({\rm X})$ is uniquely decomposed as the direct sum of the trivial representation $K$ and a possibly infinite sum of non-trivial representations, which is an ideal).
When $\rm X$ is reduced, fix-pointed means that the intersection of all non-empty closed $G$-subschemes is a $K$-rational point. In the classical situation where $K=\bar{K}$ and $G$ and $\rm X$ are of finite type over $K$, one recovers the definition in 1.1.2, cf. [10, §10].
@@ 1.3. The non-abelian Rees construction for even tensor categories
1.3.1 Theorem (O’Sullivan [36]).
For any (essentially small) even $K$-tensor category $\mathcal{C}$}, the O’Sullivan construction provides a fix-pointed affine $G$-scheme $\rm X$} and gives rise to a commutative diagram of $\otimes$-functors, the horizontal ones being equivalences.
@@
1.3.1. Example.
${\mathcal{C}}={\rm{Vec}}(\mathbb{P}^{1})$. Objects are sums of line bundles $\mathcal{O}(n)$ (Grothendieck). Here $\bar{\mathcal{C}}\cong{\rm{GrVec}}_{K},\,\bar{\omega}$ is given by the fiber at $x=\infty,\,G=\mathbb{G}_{m}$. The deformation space between $\mathcal{C}$} and $\bar{\mathcal{C}}$ is $\rm{X}=\mathbb{A}^{2}$ ($\mathbb{G}_{m}$ acting by dilatations), and one gets the diagram
Taking the fiber at the generic point of $\mathbb{P}^{1}$ embeds $\mathcal{C}$} into ${\rm{Vec}}_{K(x)}$. [^7] This goes beyond a mere exegesis of Grothendieck’s theorem on ${\rm{Vec}}(\mathbb{P}^{1})$: in fact, using his construction, O’Sullivan gave an alternative proof of Grothendieck’s theorem [37].
1.3.2. Comment on Theorem 1.3.1
(cf. also [36],[2, 3.6]). The action of $\theta$ on objects $\mathcal{V}$} of $\mathcal{C}$} (or $\bar{\mathcal{C}}$) is given by ${\mathcal{V}}\mapsto V\otimes{\mathcal{O}}({\rm X})$, where $V=\bar{\omega}({\mathcal{V}})$ as an object of ${\rm{Rep}}\,G$.
The full faithfulness of the top horizontal functor in (1.1) follows from adjunction: ${\rm{Vec}}_{G}({\rm{X}})(V\otimes{\mathcal{O}}_{\rm{X}},W\otimes{\mathcal{O}}_{\rm{X}})\cong{\rm{Hom}}_{G}(V\otimes W^{\vee},{\mathcal{O}}(\rm{X}))$
@@
1) Any object of $\mathcal{C}$} is a subobject (and also a quotient) of an object of the form $V\otimes{\mathcal{O}}_{\rm X}$ for some $V\in{\rm{Rep}}\,G$.
@@
3) If $\rm X$} is affine and ${\mathcal{O}}({\rm X})^{G}=K$ (i.e. ${\mathcal{C}}({\bf 1},{\bf 1})=K$), then the following conditions are equivalent888cf. also [10, §10]. The setting of [10] is more restrictive, but the arguments extend.:
@@
3) $a)\Rightarrow b):$ cf. [10, Cor. 6.4] (a consequence of the equivariant Nakayama lemma - the point is that $o^{\ast}$ is a conservative retraction of $p^{\ast}$ for $p:{\rm X}\to{\rm Spec}\,K$).
$b)\Rightarrow c):$ if $G^{\prime}$ is the tannakian group of $\bar{\mathcal{C}}$, there is a natural homomorphism $G^{\prime}\to G$; it induces an equivalence on representation categories, hence is an isomorphism.
$a)+c)\Rightarrow d):$ let $\rm X_{\mathcal{C}}$ be the fix-pointed affine $G$-scheme associated to $\mathcal{C}$} and the section $\sigma$. Since $\rm X$ and $\rm X_{\mathcal{C}}$ are affine, one has an essentially commutative diagram

### v2

@@ Introduction.
In the most basic case, it relates the tensor category ${\rm{FilVec}}_{K}$ of filtered $K$-vector spaces $(V,F^{\bullet})$ to the tensor category ${\rm{GrVec}}_{K}$ of graded $K$-vector spaces111here, $K$-vector spaces are finite dimensional and filtrations are decreasing, exhaustive and separated.: the map $(V,F^{\bullet})\mapsto\sum x^{-i}K[x]\otimes F^{i}V,$ (a $K[x]$-submodule of $K[x,\frac{1}{x}]\otimes V$) yields a $\mathbb{G}_{m}$-equivariant vector bundle on $\mathbb{A}^{1}={\rm{Spec}}\,K[x]$, and gives rise to a commutative diagram of tensor functors (cf. [8, §3])
@@
The present paper explores the non-abelian generalization of this situation, when $K$ is a field of characteristic 0 and the multiplicative group $\mathbb{G}_{m}$ is replaced by an arbitrary reductive $K$-group $G$. This raises immediately two complementary questions:
@@
Question 1 was given a remarkably simple and general answer by P. O’Sullivan [38]: roughly speaking, any even $K$-tensor category $\mathcal{C}$} (i.e. such that any object is killed by some exterior power) is a category of equivariant vector bundles with respect to the tannakian group $G$ of its quotient $\bar{\mathcal{C}}$ by the maximal tensor ideal, on a certain fix-pointed affine $G$-scheme $\rm X$.
Therefore, any such $\mathcal{C}$} ‘is’ the tensor category of vector bundles over some quotient stack ${\rm X}/G]$.
The question then becomes: $i)$ how to describe $\rm X$, ii)ii) how does $\rm X$ behave with respect to tensor functors ${\mathcal{C}}\to{\mathcal{C}}^{\prime}$ between even categories, and $iii)$ how properties of such functors reflect on the corresponding equivariant morphisms $\rm X^{\prime}\to\rm X$?
@@
In this paper, we first review O’Sullivan’s construction and prove its functorial properties in full generality (Theorem 1.4.2); then study in detail the case when $\mathcal{C}$} admits a faithful tensor functor with values in $K$-vector spaces. In this case, under mild extra conditions, it turns out that $\rm X$ is an affine quasi-homogeneous variety, i.e. a partial compactification of a homogeneous $G$-variety, which is amenable to computation (Theorem 2.3.1).
This gives rise to a Galois-type anti-equivalence between even categories $\mathcal{C}$} and spaces $\rm X$ (Theorem 3.2.1), which one can exploit using the rich toolbox of the theory of affine quasi-homogeneous varieties (cf. e.g. [7]).
0.6.
The non-abelian Rees construction applies notably to the setting of pure motives, especially motives modulo homological equivalence - modulo some adjustments (to cope with the Künneth projectors, see 4.2.2. The standard conjectures predict that such categories $\mathcal{C}$} of homological motives are abelian semisimple (Grothendieck’s standard conjecture D).
@@
In some special instances, the theory of affine quasi-homogeneous varieties allows to conclude that the deformation space $\rm X$ must be a point, hence that ${\mathcal{C}}=\bar{\mathcal{C}}$ is abelian semisimple. In this way, we get a new proof, and a generalization, of the Clozel-Deligne theorem on numerical equivalence on abelian varieties over finite fields (Theorem 4.6.2).
By a completely different use of the theory of affine quasi-homogeneous varieties, we also prove, over any field of characteristic $\neq\ell$, that numerical equivalence coincides with $\ell$-adic homological equivalence on powers of abelian varieties $A$ such that the center of ${\rm{End}}\,A$ is $\mathbb{Z}$ (Theorem 4.6.1).
@@ 1.1. The classical Rees construction and beyond
We are looking for extensions, when $\mathbb{G}_{m}$ is replaced by an arbitrary reductive $K$-group $G$. As mentioned above, this raises two questions Q1, Q2. In order to say a little more about Klyashko-Asok’s partial answer to question Q2, we recall three classical definitions (under the assumption $K=\bar{K}$; by ‘$K$-variety’, we mean a reduced scheme of finite type over $K$).
Quasi-homogeneous222a synonym is ‘prehomogeneous’, cf. [8]. $G$-variety: a $G$-variety $\rm{X}$} that admits a homogeneous space as an open dense subvariety ${\rm{X}_{0}}=G/H$ (informally: a partial compactification of a homogeneous space). We shall be especially concerned with such affine $G$-variety $\rm X$, which are often called ’affine embeddings of ${\rm X}_{0}$’, cf [7] for a panorama; for such affine embedding to exist, it is necessary and sufficient that $H$ is an observable subgroup of $G$.
When dealing with a quasi-homogeneous space $\rm X$, $\rm X_{0}$ will be our standard notation for its open dense homogeneous space.
Spherical $G$-variety: a quasi-homogeneous $G$-variety that remains quasi-homogeneous if one restricts the action to some Borel subgroup $B$. This condition only depends on the homogeneous space ${\rm X}_{0}$, and can be reformulated as follows: for any irreducible $G$-representation $W,\,dim\,W^{H}\leq 1$ (cf. e.g. [12]). Toric varieties are special cases, cf. e.g. [16].
Fix-pointed $G$-variety: a $G$-variety with a unique $G$-closed orbit, which is a $K$-rational point.
@@ 1.1.3. Question
It is known that every quasi-affine homogeneous variety $\rm X_{0}$ over $K=\bar{K}$ either $i)$ is affine and does not admit any non-trivial equivariant affine embedding, or ii)ii) admits an equivariant affine embedding $\rm X$ with a fixed point $o$ [7, 3.4]; but it is not clear whether $o$ may be taken to be the unique closed orbit. How restrictive is the existence of a fix-pointed equivariant affine embedding of ${\rm X}_{0}=G/H$?
1.1.4.
The Brion-Luna-Vust theory provides a combinatorial description of normal spherical varieties over $K=\bar{K}$ in terms of some $1$-parameter subgroups $\mu_{j}:\mathbb{G}_{m}\to G$ whose associated valuations333each $1$-parameter subgroup $\mu$ of $G$ defines a valuation $\nu$ on $K(G/H)$: if $f\in K[G/H]\subset K(G)$, there is a finite number of $f_{n}\in K[G]$ such that $f(g\mu(t))=\sum_{n\in\mathbb{Z}}f_{n}(g)t^{n}$ for any $g\in G(K)$, and one sets $\nu(f)={\rm{min}}\{n\mid f_{n}\neq 0\}$. on $K(G/H)$ correspond $1$-codimensional $G$-strata; in fact, for any $x\in\rm X_{0}$, ${\rm{lim}}_{t\to 0}\mu(t)x$ exists and is a point of the corresponding $1$-codimensional $G$-strata.
Using this, Asok [8, 5.7] (after Klyachko [27] in the toric case) shows that for any fix-pointed affine normal spherical $G$-variety $\rm{X}$ with open orbit ${\rm X}_{0}=G/H$, ${\rm{Vec}}_{G}({\rm X})$ is equivalent to a tensor category whose objects are $G$-representations and whose morphisms are $H$-linear maps preserving specific filtrations attached to $\mu_{j}$.
Those filtrations are defined as follows (involving the choice of auxiliary maximal tori $T_{i}$ of $G$ containing the image of $\mu_{i}$):
$F_{\mu_{i}}^{j}(V):=\oplus_{\langle\mu_{i},\chi\rangle\leq-j}\,V_{\chi},$
where $V_{\chi}$ denotes the subspace of $V\in{\rm{Rep}}\,G$ where $T_{i}$ acts through $\chi$.
@@ 1.2. O’Sullivan’s construction
This construction provides a general answer to Q1. We start with a few definitions.
@@
A $K$-tensor category $\mathcal{C}$} is a $K$-linear category444we follow O’Sullivan terminology and do not assume that $\mathcal{C}$} is abelian. equipped with a structure of symmetric monoidal category for which the tensor product is bilinear. The endomorphism ring ${{\mathcal{C}}}({\bf 1},{\bf 1})$ of the unit object is then a commutative $K$-algebra. We say that $\mathcal{C}$} is rigid if every object is dualizable.
@@
This is compatible with the usual terminology for tannakian categories, because if $\mathcal{C}$} is tannakian over $K$, any faithful tensor functor ${\mathcal{C}}\to{\rm{Vec}}_{K^{\prime}}$ is exact [40, 10.7][15, 2.4.1][24]555this result, obtained in different ways, seems to fill the gap in Saavedra’s description of tannakian categories [43] in a more economic way than Deligne’s geometry in tannakian categories [17]; cf. the discussion in [19, 3.15, p. 161], where the gap is connected to a problem of exactness..
@@ 1.2.2. Examples and notations
For a $G$-scheme $\rm X$, we denote by ${\rm{Vec}}_{G}({\rm{X}})$ the $K$-tensor category of $G$-equivariant vector bundles on $\rm X$, and write ${\rm{Vec}}({\rm{X}})$ when $G$ is trivial. Note that the tensor functor ${\rm{Vec}}({\rm{X}})\to{\rm{Vec}}_{K}$ obtained by taking the fiber at some $K$-point of $\rm X$ is not necessarily a fiber functor in our sense.
@@
A $K$-tensor category $\mathcal{C}$} is called even666or positive, in O’Sullivan’s terminology. if it is rigid and pseudo-abelian777i.e. every idempotent endomorphism has a kernel., if ${{\mathcal{C}}}({\bf 1},{\bf 1})=K$, and if every object is annihilated by some exterior power.
Then the quotient category
@@
Moreover, $\pi$ admits a tensor section $\sigma$ [38, Th. 1.1.] (the semisimplicity of $\bar{\mathcal{C}}$ is essential here), which is unique up to conjugation.
By the general criteria of SGA IV, I, 8.11, $\sigma$ then has a right Ind-adjoint
$\theta:{\rm{Ind}}\,{\mathcal{C}}\to{\rm{Ind}}\,\bar{\mathcal{C}},$
which is a lax tensor functor, and by formal reasons, $\mathcal{O}({\rm{X}}):=\theta(\bf 1)$ is a commutative algebra in ${\rm{Ind}}\,\bar{\mathcal{C}}$.
1.2.4. Reminder.
A lax tensor functor $\psi$ is defined like a tensor functor except that the natural morphisms $\psi({\mathcal{V}})\otimes\psi({\mathcal{W}})\to\psi({\mathcal{V}}\otimes{\mathcal{W}})$ are no longer supposed to be isomorphisms. If a tensor functor $\sigma:{\mathcal{B}}\to{\mathcal{A}}$ has a right adjoint $\theta$, then $\theta$ may not be a tensor functor, but it is a lax tensor functor. For any algebra object $\mathcal{V}$} in $\mathcal{A}$}, $\theta({\mathcal{V}})$ is then an algebra in $\mathcal{B}$}. A convenient setting to develop this kind of formal assertions while reducing as much as possible the diagram chases is ‘doctrinal adjunction’, in the framework of $2$-categories [29].
In this setting, the concept of mates that we shall encounter later is crucial. In the context of (essentially small) lax monoidal categories, given (strong) tensor functors
$\sigma:{\mathcal{B}}\to{\mathcal{A}},\;\sigma^{\prime}:{\mathcal{B}}^{\prime}\to{\mathcal{A}}^{\prime}$
with respective (lax tensor) right adjoints
$\theta:{\mathcal{A}}\to{\mathcal{B}},\;\theta^{\prime}:{\mathcal{A}}^{\prime}\to{\mathcal{B}}^{\prime},$
and given lax tensor functors
$\phi:{\mathcal{A}}\to{\mathcal{A}}^{\prime},\;\psi:{\mathcal{B}}\to{\mathcal{B}}^{\prime},$
one has a bijection between natural transformations of lax tensor functors
$v:\sigma^{\prime}\phi^{\prime}\Rightarrow\phi\sigma$
and natural transformations of lax tensor functors
$w:\phi^{\prime}\theta\Rightarrow\theta^{\prime}\phi$
(the mate of $v$). If $v$ is an isomorphism, $w$ may not be an isomorphism in general.
1.2.5.
In order to move toward the equivariant algebro-geometric setting, let us now assume that $\bar{\mathcal{C}}$ is neutral, and fix a fiber functor
@@
Then $G:=Aut^{\otimes}\bar{\omega}$ is a proreductive group, $\bar{\mathcal{C}}\stackrel{{\scriptstyle\sim}}{{\to}}{\rm{Rep}}\,G$, and
@@
Here fix-pointed means that there is a $G$-fixed $K$-point $o$, corresponding to a $G$-map $\mathcal{O}({\rm{X}})\to K$ whose restriction to ${\mathcal{O}}({\rm X})^{G}$ is an isomorphism. The $G$-fixed point $o$ is then unique (indeed, in ${\rm{REP}}\,G$, ${\mathcal{O}}({\rm X})$ is uniquely decomposed as the direct sum of the trivial representation $K$ and a possibly infinite sum of non-trivial representations, which is an ideal).
When $\rm X$ is reduced, ‘fix-pointed’ just means that the intersection of all non-empty closed $G$-subschemes is a $K$-rational point; and in the standard situation of $G$-varieties over $K=\bar{K}$, one recovers the definition in 1.1.2, cf. [10, §10].
@@ 1.3. The non-abelian Rees construction for even tensor categories
1.3.1 Theorem (O’Sullivan).
[38, §7] [^8] O’Sullivan works with the tensor category ${\rm Mod}(G,A)$ of dualizable $G$-AA-modules (for a $G$-algebra AA), which is equivalent to the tensor category ${\rm{Vec}}_{G}(\rm{X})$ of equivariant vector bundles over $\rm X$}= SpecA\,A. For any (essentially small) even $K$-tensor category $\mathcal{C}$} with $\bar{\mathcal{C}}$ neutral, the O’Sullivan construction provides a fix-pointed affine $G$-scheme $\rm X$} and gives rise to a commutative diagram of tensor functors, the horizontal ones being equivalences.
@@
It is not assumed that $\mathcal{C}$} admits a fiber functor.
Sketch of proof. Based on the above, the only point that remains to be seen is that the top horizontal is an equivalence.
The action of $\theta$ on objects $\mathcal{V}$} of $\mathcal{C}$} (or $\bar{\mathcal{C}}$) is given by ${\mathcal{V}}\mapsto V\otimes{\mathcal{O}}({\rm X})$, where $V=\bar{\omega}({\mathcal{V}})$ as an object of ${\rm{Rep}}\,G$.
Full faithfulness of the top horizontal functor follows from adjunction:
${\rm{Vec}}_{G}({\rm{X}})(V\otimes{\mathcal{O}}_{\rm{X}},W\otimes{\mathcal{O}}_{\rm{X}})\cong{\rm{Hom}}_{G}(V\otimes W^{\vee},{\mathcal{O}}(\rm{X}))$
@@
1) Any object of $\mathcal{C}$} is a quotient (and also a subobject) of an object of the form $V\otimes{\mathcal{O}}_{\rm X}$ for some $V\in{\rm{Rep}}\,G$.
@@
3) If $\rm X$} is affine and ${\mathcal{O}}({\rm X})^{G}=K$ (i.e. ${\mathcal{C}}({\bf 1},{\bf 1})=K$), then the following conditions are equivalent999cf. also [10, §10]. The setting of [10] is more restrictive, but the arguments extend.:
@@
3) $a)\Rightarrow b):$ cf. [10, Cor. 6.4] (a consequence of the equivariant Nakayama lemma - the point is that $o^{\ast}$ is a conservative retraction of $p^{\ast}$ for the structural map $p:{\rm X}\to{\rm Spec}\,K$).
$b)\Rightarrow c):$ if $G^{\prime}$ is the tannakian group of $\bar{\mathcal{C}}$, there is a natural homomorphism $G^{\prime}\to G$ of proreductive groups. At the level of (semisimple) representation categories, it induces a bijection on isomorphism classes of objects, hence is a $\otimes$-equivalence. Therefore $G^{\prime}\to G$ is an isomorphism.
$a)+c)\Rightarrow d):$ let $\rm X_{\mathcal{C}}$ be the fix-pointed affine $G$-scheme associated to $\mathcal{C}$} and the section $\sigma$ by O’Sullivan’s construction. Since $\rm X$ and $\rm X_{\mathcal{C}}$ are affine, one has an essentially commutative diagram
@@
1.3.1. Example.
${\mathcal{C}}={\rm{Vec}}(\mathbb{P}^{1})$. Objects are sums of line bundles $\mathcal{O}(n)$ (Grothendieck), and $\mathcal{C}$} is $\otimes$-equivalent to ${\rm{Vec}}_{\mathbb{G}_{m}}(\mathbb{A}^{2}\setminus\{0\})$ ($\mathbb{G}_{m}$ acting by dilatations), hence also to ${\rm{Vec}}_{\mathbb{G}_{m}}(\mathbb{A}^{2})$ since we are in dimension $2$. The deformation space between $\mathcal{C}$} and $\bar{\mathcal{C}}$ is $\rm{X}=\mathbb{A}^{2}$, and one gets the diagram
Taking the fiber at the generic point of $\mathbb{P}^{1}$ embeds $\mathcal{C}$} into ${\rm{Vec}}_{K(x)}$. This goes beyond a mere exegesis of Grothendieck’s theorem on ${\rm{Vec}}(\mathbb{P}^{1})$: in fact, using his construction, O’Sullivan gave an alternative proof of Grothendieck’s theorem [39].